OpenStudy (mendicant_bias):

Dealing with a problem involving Stokes Theorem with a vector field in R3 and a line integral in R2; a little confused about something that pops up because of it.

OpenStudy (mendicant_bias):

Posted below in a moment, just one second. http://i.imgur.com/HMz7sqa.png

OpenStudy (mendicant_bias):

What should I do about that z in the integrand/is that okay?

ganeshie8 (ganeshie8):

z = 0 if you're living in xy plane

OpenStudy (mendicant_bias):

So does the integrand just become -1? e.g. in any situation where I took the Curl of a vector field and got some result where there was a nonzero component parallel to my normal, and \[\triangledown \times F \cdot n \neq 0 \](Lol redundant)-

OpenStudy (mendicant_bias):

I can just eliminate it or plug in the value that it constantly is?

ganeshie8 (ganeshie8):

Yes which ever way u prefer

OpenStudy (mendicant_bias):

Alright, so if I had a situation where my normal vector was, say, \[j\] and my Curl vector was \[x i + 12y \ j - 3k,\]The dot product and thus integrand would be 12y, and I could *not* progress or go further in the integral without simplifying, or I could just factor it out and treat it like a constant? (Now continuing with this problem)

OpenStudy (mendicant_bias):

\[\int\limits_{-3}^{3}\int\limits_{-\sqrt{9-y^2}}^{\sqrt{9-y^2}}(-1) \ dx \ dy \]Is this general setup correct for cartesian?

ganeshie8 (ganeshie8):

`12y` is not a constant so you need to setup bounds and evaluate the integral the usual way

ganeshie8 (ganeshie8):

yes but "-1" is constant here, simply multiply the area of circle by the curl

OpenStudy (mendicant_bias):

Well I mean, if your line integral was lying in the xz plane and thus the normal unit vector was j hat, could you treat y like a constant if you got y in your grad cross F dot n, or even factor it out like a constant?

ganeshie8 (ganeshie8):

y is 0 if your path is in xz plane

OpenStudy (mendicant_bias):

(Will do, just talking about that last part first)-Oh, or parallel to the xz plane, just any plane parallel or including the xz plane, could you treat it as a constant. I know I can do this, heh, but I'm really unfamiliar with it here; so when I have this area element dA, how do I systematically do the thing you're talking about, multiplying by the area of the object?

ganeshie8 (ganeshie8):

\[ \large \begin{align} \int\limits_{-3}^{3}\int\limits_{-\sqrt{9-y^2}}^{\sqrt{9-y^2}}(-1) \ dx \ dy &= -\int\limits_{-3}^{3}\int\limits_{-\sqrt{9-y^2}}^{\sqrt{9-y^2}} 1 \ dx \ dy \\~\\ &= - (\text{Area of circle of radius 3})\\~\\ &= -(\pi\cdot 3^2) \end{align} \]

ganeshie8 (ganeshie8):

Is that what you talking about ?

OpenStudy (mendicant_bias):

Yeah, no, that makes sense, I'm just being dumb and not thinking in terms of what integration actually is. Lemme check to see if I got the right answer.

OpenStudy (mendicant_bias):

My book says it's supposed to be \[-\frac{5}{6}\]

OpenStudy (mendicant_bias):

Going to look for a mistake, but I don't yet see where I made one; no transformation this time.

OpenStudy (mendicant_bias):

GUHHHH >:C I MISREAD THE PROBLEM, the last term in the vector field is +x^2, not -z^2

ganeshie8 (ganeshie8):

it wont change the answer because we don't care about z component

OpenStudy (mendicant_bias):

It will change the Curl vector, though, won't it? It will make some things not disappear and some other things appear.

OpenStudy (mendicant_bias):

We now have a nonzero j component in the Curl vector, if I think right

ganeshie8 (ganeshie8):

try it, you will get the same value for curl

ganeshie8 (ganeshie8):

j component or k component ?

ganeshie8 (ganeshie8):

you're changing k component from -z^2 to x^2 right ?

OpenStudy (mendicant_bias):

J component, since it's a determinant, it changes a good bit :E Yeah

OpenStudy (mendicant_bias):

Lemme write out and show my setup comparing the two.

ganeshie8 (ganeshie8):

how does changing z component of force affect the things in xy plane ?

ganeshie8 (ganeshie8):

you will get the same "-1" for integrand, work it and see

OpenStudy (mendicant_bias):

I see what you mean, but yeah, it does change the curl vector, which is what I was saying, heh. http://i.imgur.com/ZrNImYc.png Now I just need to find my mistakes otherwise.

ganeshie8 (ganeshie8):

yeah i see :O

OpenStudy (mendicant_bias):

Going to triple check that I read both the problem and the answer key correctly.

OpenStudy (mendicant_bias):

>:CCCCCCCCCCCC

ganeshie8 (ganeshie8):

lol

OpenStudy (mendicant_bias):

I switched an entire problem, I did number two, which there is no answer for in the answer key! I'll close this one and do it entirely separately, MAN

ganeshie8 (ganeshie8):

:/

OpenStudy (mendicant_bias):

Lul, my second to last comment made it light-hearted enough for me. Opening up the new question in a minute. It might be really similar IIRC, so I may not even open the new one, will see momentarily.

ganeshie8 (ganeshie8):

ok